OpenStudy (anonymous):

Do 56, 90, and 106 form a pythgorean triple? Why or why not?

OpenStudy (anonymous):

yes, 56^2+90^2=106^2

OpenStudy (anonymous):

a^2 +b^2=c^2

OpenStudy (anonymous):

Do you know what it means, Pythagorean triple?

OpenStudy (anonymous):

Yeah.. I was just confused for second.. But thanks